OpenStudy (anonymous):

In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10 1/3x + 2 -1/3x + 6 –x + 2 -1/3x + 2

ganeshie8 (ganeshie8):

take the first equation, and try to isolate y

ganeshie8 (ganeshie8):

x – 3y = –6

OpenStudy (anonymous):

how do i do that? you're gonna have to take baby steps haha

ganeshie8 (ganeshie8):

aha..

ganeshie8 (ganeshie8):

first baby step : subtract x both sides

OpenStudy (anonymous):

alrighty...

ganeshie8 (ganeshie8):

walk a bit, and show me wat u get

OpenStudy (anonymous):

i really have no idea i'm sorry haha

ganeshie8 (ganeshie8):

's okay :)

ganeshie8 (ganeshie8):

x – 3y = –6 subtract x both sides x – 3y = –6 -x -x

OpenStudy (anonymous):

alright...

ganeshie8 (ganeshie8):

-3y = -x-6

OpenStudy (anonymous):

oh thats what i just got

ganeshie8 (ganeshie8):

good, next see that -3 is attached to y

ganeshie8 (ganeshie8):

since we want to solve y, we need to get rid of it

ganeshie8 (ganeshie8):

just we do the opposite, divide both sides by -3

ganeshie8 (ganeshie8):

-3y = -x-6 /-3 /-3

ganeshie8 (ganeshie8):

y = 1/3x + 2

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

oh that does make sense

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

so it's a right?

ganeshie8 (ganeshie8):

thats right, its A

OpenStudy (anonymous):

awesome that's what i got when i tried it out but i wasnt positive, thank you!

ganeshie8 (ganeshie8):

np :) you're wlcme !!